### v2

# Extremal metrics involving scalar curvature

Abstract
We investigate extremal metrics at which various types of rigidity theorems involving scalar curvatures hold. The rigidity we discuss here is related to the rigidity theorems presented by Mario Listing in his previous preprint. More specifically, we give some sufficient conditions for metrics not to be rigid in this sense. We also give several examples of Riemannian manifolds that satisfy such sufficient conditions.
footnotetext: Keywords: Scalar curvature rigidityfootnotetext: 2020 Mathematics Subject Classification: 53C20, 53C24
1 Introduction
Llarull [21] showed some rigidity results for the standard sphere. And, Goette and Semmelmann [10] generalized it to locally symmetric spaces of compact type and nontrivial Euler characteristic. Later Listing [19] generalized their results in the following form.
Theorem 1.1 ([19, Theorem 1]).
Let $(M_{0}^{n},g_{0})~{}(n\geq 3)$ be an oriented closed Riemannian manifold with nonnegative curvature operator, positive Ricci curvature and non-vanishing Euler characteristic. Suppose that $(M^{n},g)$ is an oriented closed Riemannian manifold and $f:M\rightarrow M_{0}$ is a spin map of non-zero degree. If the scalar curvature satisfies
$R_{g}\geq(R_{g_{0}}\circ f)\cdot\sqrt{\mathrm{area}(f)},$
then $\alpha:=\mathrm{area}(f)$ is a (positive) constant and $f:(M,\alpha\cdot g)\rightarrow(M_{0},g_{0})$ is a Riemannian covering. Here, $R_{g},R_{g_{0}}$ denote the scalar curvature of $g,g_{0}$ respectively and
$\mathrm{area}(f):M\rightarrow[0,\infty);~{}x\mapsto\max_{v\in\Lambda^{2}T_{x}M\setminus\{0\}}\frac{f^{*}g_{0}(v,v)}{g(v,v)}.$
For the case that $M=M_{0}$ and $f=\mathrm{id}_{M}$, he also gave the following type of rigidity theorem.
Theorem 1.2 ([19, Theorem 2]).
Suppose $(M^{n},g_{0})$ is an oriented Riemannian manifold of dimension $n=4k+1~{}(k\in\mathbb{N})$ with nonnegative curvature operator, positive Ricci curvature and non-vanishing Kervaire semicharacteristic $\sigma(M)\neq 0$. If $g$ is a Riemannian metric on $M$ satisfying
$R_{g}\geq R_{g_{0}}\cdot\|g_{0}\|_{2,g},$
then there is a positive constant $c>0$ such that $g=c\cdot g_{0}$. Here, $\|g_{0}\|_{2,g}=\mathrm{area}(\mathrm{id}_{M})$ is defined by (6) below.
Let $(M,g_{0})$ be a smooth Riemannian manifold and $\mathcal{M}$ the space of all Riemannian metrics on $M$. Consider the functional
$R_{min}:\mathcal{M}\ni g\mapsto\min_{M}R_{g}\in\mathbb{R}$
and a functional $\mu_{g_{0}}$ on $\mathcal{M}$, which is determined by $g_{0}$ and scaling invariant of weight $-1$, i.e., $\mu(c\cdot g)=c^{-1}\mu(g)$ for all $c>0$ and $g\in\mathcal{M}$. If the metric $g_{0}$ is rigid with respect to the functional $\mu_{g_{0}}$ in a certain sense, then $\mu_{g_{0}}$ is an upper bound of $R_{min}$ as a functional on $\mathcal{M}$ and these values coincide at $g_{0}$. On the other hand, when $M$ is closed $n$-manifold with non-positive Yamabe invariant $Y(M)$, the following functional is an upper bound of $R_{min}$ on $\mathcal{M}$:
$g\mapsto Y(M)\cdot\mathrm{Vol}(M,g)^{-2/n}.$
Moreover, if a smooth metric $g_{0}$ attains the equality, then it is a Yamabe metric (i.e., $Y(M,[g_{0}])=Y(M)$) and an Einstein metric. However, when the Yamabe invariant is positive, this is not the case in general (see Remark 2.1 below or [20, Chapter 3]). A Yamabe metric is expected to be standard in some sense (cf. [28, Section 1]), but it remains a difficult problem to know how to actually obtain it as a limit of the sequence of solutions to the Yamabe problem, and in what sense it is standard (see also Subsection 6.2 below).
As noted in [10], one can apply the construction of Lohkamp [22] to see that not all metrics on $M$ are area-extremal (for this definition, see the beginning of Section 2) if $\mathrm{dim}M\geq 3$. However, such an example is not given in an explicit way. That is, we can deduce the existence of such a metric but we cannot know any concrete properties involving its curvatures in general. In light of the above, this study aims to investigate relations between some rigidity phenomena involving scalar curvature and standard metrics in various senses. In particular, in this paper, we give some necessary conditions for metrics to be extremal in some senses (see Definitions 1.1, 1.2 and 1.3 below and the corollaries following them).
Throughout the paper, any Riemannian metric will be smooth. The Ricci curvature and the scalar curvature of a metric $g$ are denoted by $\mathrm{Ric}_{g}$ and $R_{g}$, respectively. The (non-positive) Laplacian that acts on functions is defined by $\Delta_{g}f=\mathrm{tr}_{g}\nabla^{g}df$, where $\nabla^{g}$ is the Levi-Civita connection of $g$. The volume element is denoted by $\mathrm{vol}_{g}$. For two symmetric $(0,2)$-tensors $g$ and $h$, we say $g\geq h$ on $\Lambda^{2}TM$ if $g(v,w)\geq h(v,w)$ for all $v,w\in\Lambda^{2}TM$.
Our first main theorem is the following.
Theorem 1.3.
Let $(M^{n},g)$ be a closed Riemannian manifold of dimension $n\geq 2$ satisfying that
$\begin{split}&~{}~{}~{}\frac{3n-2}{2n}\Delta_{g}R_{g}+\|\overset{\circ}{\mathrm{Ric}_{g}}\|^{2}_{g}+\frac{R_{g}^{2}}{n}>R_{g}\cdot\max_{v\in T_{x}M,\,|v|_{g}=1}\mathrm{Ric}_{g}(v,v)~{}~{}\mathrm{on}~{}M,\\ &\left(\mathrm{resp.}\,\frac{3n-2}{2n}\Delta_{g}R_{g}+\|\overset{\circ}{\mathrm{Ric}_{g}}\|^{2}_{g}+\frac{R_{g}^{2}}{n}>R_{g}\cdot\min_{v\in T_{x}M,\,|v|_{g}=1}\mathrm{Ric}_{g}(v,v)~{}~{}\mathrm{on}~{}M\right)\end{split}$ (1)
where $\overset{\circ}{\mathrm{Ric}_{g}}:=\mathrm{Ric}_{g}-\frac{R_{g}}{n}g$ is the traceless Ricci tensor. Then, there is a small constant $s>0$ (resp. $s<0$) depending only on $n,M$ and $g$ such that $g_{s}=g-s\,\overset{\circ}{\mathrm{Ric}_{g}}$ is a Riemmanian metric on $M$ and that
$R_{g_{s}}>R_{g}\cdot\|g\|^{2}_{1,g_{s}}~{}~{}(\mathrm{resp.}~{}R_{g_{s}}<R_{g}\cdot\|g\|^{2}_{1,g_{s}})$ (2)
at each point of $M$. Here, $\|g\|_{1,g_{s}}:M\rightarrow[0,\infty)$ is the function on $M$ defined by
$\|g\|_{1,g_{s}}(x):=\sqrt{\max_{v\in T_{x}M\setminus\{0\}}\frac{g(v,v)}{g_{s}(v,v)}}.$ (3)
Remark 1.1.
In particular, if $(M,g)$ has a negative constant scalar curvature and satisfies
$\|\overset{\circ}{\mathrm{Ric}_{g}}\|_{g}(x)\neq 0~{}~{}\mathrm{for~{}all}~{}x\in M,$ (4)
then $(M,g)$ satisfies the assumption in the first line of (1). On the other hand, if $(M,g)$ has a positive constant scalar curvature and satisfies (4), then it satisfies the assumption in the second line of (1).
Similarly to Theorem 1.3, we can also prove the following.
Theorem 1.4.
Let $(M^{n},g)$ be a closed Riemannian manifold of dimension $n\geq 2$ satisfying the following. Let $\lambda_{Ric,1}(x)\leq\cdots\leq\lambda_{Ric,n-1}(x)\leq\lambda_{Ric,n}(x)$ be the ordered eigenvalues of $\mathrm{Ric}_{g}$ on $T_{x}M~{}(x\in M)$.
  • $R_{g}\leq 0$ on $M$ and on $M$,
    $\begin{split}&~{}~{}~{}~{}\|\overset{\circ}{\mathrm{Ric}_{g}}\|_{g}^{2}+\frac{3n-2}{2n}\Delta_{g}R_{g}+\frac{R_{g}^{2}}{n}-R_{g}\cdot\lambda_{Ric,n-1}>0\\ &\left(\mathrm{resp.}~{}\|\overset{\circ}{\mathrm{Ric}_{g}}\|_{g}^{2}+\frac{3n-2}{2n}\Delta_{g}R_{g}+\frac{R_{g}^{2}}{n}-R_{g}\cdot\lambda_{Ric,2}>0\right)\end{split}$
    or
  • $R_{g}\geq 0$ on $M$ and on $M$,
    $\begin{split}&~{}~{}~{}~{}\|\overset{\circ}{\mathrm{Ric}_{g}}\|_{g}^{2}+\frac{3n-2}{2n}\Delta_{g}R_{g}+\frac{R_{g}^{2}}{n}-R_{g}\cdot\lambda_{Ric,n}>0\\ &\left(\mathrm{resp.}~{}\|\overset{\circ}{\mathrm{Ric}_{g}}\|_{g}^{2}+\frac{3n-2}{2n}\Delta_{g}R_{g}+\frac{R_{g}^{2}}{n}-R_{g}\cdot\lambda_{Ric,1}>0\right)\end{split}$
Then, there is a small constant $s>0$ (resp. $s<0$) depending only on $n,M$ and $g$ such that $g_{s}=g-s\,\overset{\circ}{\mathrm{Ric}_{g}}$ is a Riemmanian metric on $M$ and that
$R_{g_{s}}>R_{g}\cdot\|g\|_{2,g_{s}}~{}~{}(\mathrm{resp.}~{}R_{g_{s}}<R_{g}\cdot\|g\|_{2,g_{s}})$ (5)
at each point of $M$. Here, $\|g\|_{2,g_{s}}:M\rightarrow[0,\infty)$ is the function on $M$ defined by
$\|g\|_{2,g_{s}}(x):=\sqrt{\max_{v\in\Lambda^{2}T_{x}M\setminus\{0\}}\frac{g(v,v)}{g_{s}(v,v)}}.$ (6)
Theorem 1.5.
Let $(M^{n},g)$ be a closed Riemannian manifold of dimension $n\geq 2$ satisfying that
$\left(\frac{3n-2}{2n}\Delta_{g}R_{g}+\|\overset{\circ}{\mathrm{Ric}_{g}}\|^{2}_{g}+\frac{R_{g}^{2}}{n}\right)\cdot g>R_{g}\cdot\mathrm{Ric}_{g}~{}\mathrm{on}~{}TM.$
Then, there is a small constant $s>0$ (resp. $s<0$) depending only on $n,M$ and $g$ such that $g_{s}=g-s\,\overset{\circ}{\mathrm{Ric}_{g}}$ is a Riemmanian metric on $M$ and that
$R_{g_{s}}\cdot g_{s}>R_{g}\cdot g~{}~{}(\mathrm{resp.}~{}R_{g_{s}}\cdot g_{s}<R_{g}\cdot g)$ (7)
at each point of $M$.
Theorem 1.6.
Let $(M^{n},g)$ be a closed Riemannian manifold of dimension $n\geq 2$ satisfying that
$\left(\frac{3n-2}{n}R_{g}\,\Delta_{g}R_{g}+2R_{g}\,\|\overset{\circ}{\mathrm{Ric}_{g}}\|^{2}_{g}+\frac{R_{g}^{3}}{n}\right)\cdot g>R^{2}_{g}\cdot\mathrm{Ric}_{g}~{}\mathrm{on}~{}\Lambda^{2}TM.$
Then, there is a small constant $s>0$ (resp. $s<0$) depending only on $n,M$ and $g$ such that $g_{s}=g-s\,\overset{\circ}{\mathrm{Ric}_{g}}$ is a Riemmanian metric on $M$ and that
$R^{2}_{g_{s}}\cdot g_{s}>R^{2}_{g}\cdot g~{}~{}(\mathrm{resp.}~{}R^{2}_{g_{s}}\cdot g_{s}<R^{2}_{g}\cdot g)$ (8)
on $\Lambda^{2}TM$.
Remark 1.2.
  • All the assumptions of the above Theorems 1.3, 1.4, 1.5 and 1.6 especially imply that
    $\|\overset{\circ}{\mathrm{Ric}_{g}}\|_{g}\neq 0~{}~{}\mathrm{on}~{}M.$
  • If $\Delta_{g}R_{g}=0$ on $M$, then $R_{g}$ is sign-changing otherwise $R_{g}\equiv\mathrm{const}$ on $M$.
Interestingly, Dahl–Kröncke [7, 16] recently also discovered a relation between stability of Einstein metrics and certain type of scalar curvature rigidity. Since every Einstein metric does not satisfy any of the above assumptions (see Remark 1.2 above), our theorems above cannot be applied to Einstein metrics. However, we can also prove the following.
Theorem 1.7.
Let $(M^{n},g)$ be a closed non-positive Einstein manifold of dimension $n\geq 2$ with unit volume. Assume either that
  • $(M,g)$ is Ricci flat, unstable ([16]), i.e., there is a tt-tensor $h$ (see 4 for the definition) such that
[…]
    where $\Delta_{E}:=\nabla_{g}^{*}\nabla_{g}-2\overset{\circ}{R}$ is the Einstein tensor and
    $(\overset{\circ}{R}h)(X,Y):=\sum_{i}h(R(X,e_{i})Y,e_{i}).$
    Moreover, there is a corresponding curve of metrics $g_{s}~{}(|s|<\varepsilon)$ lying in the space of all Riemannian metrics of unit volume with constant scalar curvature such that $\left.\frac{d}{ds}g_{s}\right|_{s=0}=h$ and $\left.\frac{d^{2}}{d^{2}s}g_{s}\right|_{s=0}$ is also a tt-tensor,
  • $R_{g}<0$ and there is a non-vanishing tt-tensor $h$, i.e.,
    $\max_{M}\sup\{C\in\mathbb{R}\mid~{}h_{x}(v,v)\geq C\cdot g_{x}(v,v)~{}\mathrm{for~{}all}~{}v\in T_{x}M\}<0.$
Then, there is a curve of metrics $g_{s}$ defined for sufficiently small $s>0$ with $\frac{d}{ds}g_{s}|_{s=0}=h$ such that $g_{s}$ satisfies
$R_{g_{s}}>R_{g}\cdot\|g\|_{1,g_{s}}^{2}$ (10)
on $M$. Here, $\|g\|_{1,g_{s}}$ is the function defined in (3).
Remark 1.3.
All known examples of unstable Einstein manifolds are of positive scalar curvature (see [15]). Therefore, it is not known whether there is a metric that satisfies the first condition.
Question 1.1 ([8, p. 65]).
Are all compact Einstein manifolds with non-positive scalar curvature stable?
Theorem 1.8.
Let $(M^{n},g)$ be a closed Einstein manifold of dimension $n\geq 2$ with unit volume. Assume that $(M,g)$ is not isometric to the standard round sphere. For a tt-tensor $h$ (see 4 for the definition), we denote the second eigenvalue of $h$ with respect to $g$ as $\lambda^{h}_{2}$. Assume either that
  • $R_{g}\neq 0$ and there is a tt-tensor $h$ such that $\max_{M}\lambda^{h}_{2}<0$, or
  • $R_{g}=0$, $g$ is unstable and a tt-tensor $h$ satisfying (9). Moreover, there is a corresponding curve of metrics $g_{s}~{}(|s|<\varepsilon)$ lying in the space of all Riemannian metrics of unit volume with constant scalar curvature such that $\left.\frac{d}{ds}g_{s}\right|_{s=0}=h$ and $\left.\frac{d^{2}}{d^{2}s}g_{s}\right|_{s=0}$ is also a tt-tensor.
Then, there is a curve of metrics $g_{s}~{}(|s|<\varepsilon)$ such that for all $s\in(0,\varepsilon)$, $g_{s}$ satisfies
$R_{g_{s}}>R_{g}\cdot\|g\|_{2,g_{s}}$
on $M$. Here, $\|g\|_{2,g_{s}}$ is the function defined in (6).
Remark 1.4.
In the case of $R_{g}\neq 0$, the assumption “$\max_{M}\lambda^{h}_{2}<0$” especially implies that $h$ is non-vanishing on $M$.
According to Listing’s work [19], we define three types of rigidities of metrics involving scalar curvature.
Definition 1.1.
Let $M$ be a smooth Riemannian manifold. A metric $g_{0}$ is type I scalar curvature rigid in the sense of Listing if for any metric $g$ on $M$,
$R_{g}\geq R_{g_{0}}\cdot\|g_{0}\|^{2}_{1,g}$ (11)
implies that $g=c\cdot g_{0}$ for some positive constant $c>0$. Here, $\|g\|_{1,g_{s}}$ is the function defined in (3).
Definition 1.2.
Let $M$ be a smooth Riemannian manifold. A metric $g_{0}$ is type II scalar curvature rigid in the sense of Listing if for any metric $g$ on $M$,
$R_{g}\geq R_{g_{0}}\cdot\|g_{0}\|_{2,g},$ (12)
implies that $g=c\cdot g_{0}$ for some positive constant $c>0$. Here, $\|g\|_{2,g_{s}}$ is the function defined in (6).
Definition 1.3.
Let $M$ be a smooth Riemannian manifold. A metric $g_{0}$ is type III scalar curvature rigid in the sense of Listing if for any metric $g$ on $M$,
$R_{g}\cdot g\geq R_{g_{0}}\cdot g_{0}~{}~{}\mathrm{on}~{}TM,$ (13)
implies that $g=c\cdot g_{0}$ for some positive constant $c>0$.
Definition 1.4.
Let $M$ be a smooth Riemannian manifold. A metric $g_{0}$ is type IV scalar curvature rigid in the sense of Listing if for any metric $g$ on $M$,
$R_{g}^{2}\cdot g\geq R_{g_{0}}^{2}\cdot g_{0}~{}~{}\mathrm{on}~{}\Lambda^{2}TM,$ (14)
implies that $g=c\cdot g_{0}$ for some positive constant $c>0$.
Remark 1.5.
The condition (13) implies (11). And, if $R_{g_{0}}\geq 0$ on $M$, then the condition (12) is equivalent to (14).
By taking the contrapositive of each of the above theorems, we obtain necessary conditions for metrics to be scalar curvature rigid in the sense of the above definitions.
Corollary 1.1.
Let $M^{n}$ be a closed manifold of dimension $n\geq 2$. If a metric $g$ on $M$ is type I scalar curvature rigid in the sense of Listing, then
$\frac{3n-2}{2n}\Delta_{g}R_{g}(x)+\|\overset{\circ}{\mathrm{Ric}_{g}}\|_{g}^{2}(x)+\frac{R_{g}^{2}(x)}{n}\leq R_{g}(x)\cdot\max_{v\in T_{x}M,\,|v|_{g}=1}\mathrm{Ric}_{g}(v,v)$
at some point $x\in M$.
Corollary 1.2.
Let $M^{n}$ be a closed manifold of dimension $n\geq 2$. If a metric $g$ on $M$ is type II scalar curvature rigid in the sense of Listing, then either of the following holds:
  • $\max_{M}R_{g}>0$ or
    $\frac{3n-2}{2n}\Delta_{g}R_{g}(x)+\|\overset{\circ}{\mathrm{Ric}_{g}}\|_{g}^{2}(x)+\frac{R_{g}^{2}(x)}{n}-R_{g}(x)\cdot\lambda_{Ric,n-1}(x)\leq 0$
    at some point $x\in M$, and
  • $\min_{M}R_{g}<0$ or
    $\frac{3n-2}{2n}\Delta_{g}R_{g}(x)+\|\overset{\circ}{\mathrm{Ric}_{g}}\|_{g}^{2}(x)+\frac{R_{g}^{2}(x)}{n}-R_{g}(x)\cdot\lambda_{Ric,n}(x)\leq 0$
    at some point $x\in M$.
Corollary 1.3.
Let $M^{n}$ be a closed manifold of dimension $n\geq 2$. If a metric $g$ on $M$ is type III scalar curvature rigid in the sense of Listing, then
$\left(\frac{3n-2}{n}\Delta_{g}R_{g}(x)+2\|\overset{\circ}{\mathrm{Ric}_{g}}\|^{2}_{g}(x)+\frac{R_{g}^{2}(x)}{n}\right)\cdot g_{x}(v,w)\leq R_{g}(x)\cdot\mathrm{Ric}_{g}(x)(v,w)$
at some point $x\in M$ and some vectors $v,w\in T_{x}M$.
Corollary 1.4.
Let $M^{n}$ be a closed manifold of dimension $n\geq 2$. If a metric $g$ on $M$ is type IV scalar curvature rigid in the sense of Listing, then
$\begin{split}&\left(\frac{3n-2}{n}R_{g}(x)\,\Delta_{g}R_{g}(x)+2R_{g}(x)\,\|\overset{\circ}{\mathrm{Ric}_{g}}\|^{2}_{g}(x)+\frac{R_{g}^{3}(x)}{n}\right)\cdot g_{x}(v,w)\\ &\leq R^{2}_{g}(x)\cdot\mathrm{Ric}_{g}(x)(v,w)\end{split}$
at some point $x\in M$ and some two-vectors $v,w\in\Lambda^{2}T_{x}M$.
Corollary 1.5.
Let $(M^{n},g)$ be a closed negative Einstein manifold of dimension $n\geq 2$ with unit volume. If it is type I scalar curvature rigid in the sense of Listing, then $h_{x}=0$ at some point $x\in M$ for any tt-tensor $h$.
Corollary 1.6.
Let $(M^{n},g)$ be a closed non-Ricci-flat Einstein manifold of dimension $n\geq 2$ with unit volume that is not isometric to the standard sphere. If it is type II scalar curvature rigid in the sense of Listing, then $\min_{M}\lambda^{h}_{2}\geq 0$ for any tt-tensor $h$. Here, $\lambda^{h}_{2}$ denote the second eigenvalue of $h$ with respect to $g$.
This paper is organized as follows. In Section 2, we review some rigidity results for certain Riemannian metrics with positive scalar curvature and a relation between “scalar minimum functional” and an extremal metric (see Remark 2.1). In Section 3, we describe a formula that is necessary to prove our main theorems. Furthermore, we consider statements of the same type as our main theorems on compact manifolds with boundary and non-compact complete manifolds. In Section 4, we prove our main theorems. In Section 5, we give some examples that satisfy the assumptions of Theorem 1.3, 1.4, 1.5 and 1.6. In Section 6, we present some further questions related to our main theorems. In Section 7, we give a proof of the formula in Section 3.
2 Previous rigidity results for metrics with positive scalar curvature
A metric $g$ on a smooth manifold $M$ is called (globally) area-extremal if, for a metric $h$ satisfying $h\geq g$ on $\Lambda^{2}TM$, $R_{h}\geq R_{g}$ holds only when $R_{h}=R_{g}$ on $M$. As a generalization of Llarull’s significant rigidity result [21], Goette and Semmelmann [10] gave a sufficient condition for a metric to be locally area-extremal as follows.
Proposition 2.1 ([10, Lemma 0.3]).
Let $(M,g)$ be a compact Riemannian manifold whose Ricci curvature $\mathrm{Ric}_{g}$ is positive definite on $M$. Then there exists no nonconstant $C^{1}$-path $(g_{t})_{t\in[0,\varepsilon]}$ of Riemannian metrics on $M$ for $\varepsilon>0$ with $g_{0}=g$, such that $g_{t}\geq g$ on $TM$ and $R_{g_{t}}\geq R_{g_{0}}$ on $M$.
Suppose moreover that $2\,\mathrm{Ric}_{g}-R_{g}\cdot g$ is negative definite on $M$. Then there is no nonconstant path $(g_{t})_{t\in[0,\varepsilon]}$ as above, such that $g_{t}\geq g$ on $\Lambda^{2}TM$ and $R_{g_{t}}\geq R_{g}$ on $M$.
Meanwhile, they also gave the following stability result.
Theorem 2.1 ([10, Theorem 2.4]).
Let $(M_{0}^{n},g_{0})~{}(n\geq 3)$ be an oriented closed Riemannian manifold with nonnegative curvature operator, positive Ricci curvature and non-vanishing Euler characteristic. Suppose that $(M^{n},g)$ is an oriented closed Riemannian manifold and $f:M\rightarrow M_{0}$ is a spin map of non-vanishing $\hat{A}$-degree $\mathrm{deg}_{\hat{A}}(f)\neq 0$ and $\mathrm{area}(f)\leq 1$. Then $R_{g}\geq R_{g_{0}}\circ f$ implies that $R_{g}=R_{g_{0}}\circ f$. If moreover, $\mathrm{Ric}_{g}>0$ and $02\,\mathrm{Ric}_{g}-R_{g}\cdot g<$ 0 on $M$, then $f:M\rightarrow M_{0}$ is a Riemannian submersion.
They also prove area-extremality and rigidity for a certain class of metrics with nonnegative curvature operator on $\Lambda^{2}TM$.
Theorem 2.2 ([10]).
Let $(M,g)$ be a compact connected oriented Riemanniam manifold with nonnegatuve curvature operator on $\Lambda^{2}TM$, such that the universal covering of $M$ is homeomorphic to a symmetric space $G/K$ of compact type with $\mathrm{rk}G\leq\mathrm{rk}K+1$. Then $g$ is (globally) area-extremal. If moreover, $\mathrm{Ric}_{g}>0$ and $02\,\mathrm{Ric}_{g}-R_{g}\cdot g<$ 0 on $M$, then $R_{h}\geq R_{g}$ and $h\geq g$ on $\Lambda^{2}TM$ implies $h=g$.
Later Listing generalized these to Theorem 1.1 and 1.2 above. On the other hand, Lott [23] extended results of Llarull and Goette–Semmelmann to manifolds with boundary.
Remark 2.1.
Let $M^{n}$ be a closed manifold of dimension $n\geq 2$ and $\mathcal{M}(M)$ the space of all Riemannian metrics on $M$. Consider the following scalar minimum functional:
$R_{min}:\mathcal{M}(M)\rightarrow\mathbb{R}~{};~{}g\mapsto\min_{M}R_{g}.$
For a fixed Riemannian metric $g_{0}\in\mathcal{M}(M)$, we define the following two functionals $F_{1}(g_{0})$ and $F_{2}(g_{0})$.
$\begin{split}F_{1,g_{0}}&:\mathcal{M}(M)\rightarrow\mathbb{R}~{};~{}g\mapsto R_{g_{0}}\cdot\|g_{0}\|^{2}_{1,g},\\ F_{2,g_{0}}&:\mathcal{M}(M)\rightarrow\mathbb{R}~{};~{}g\mapsto R_{g_{0}}\cdot\|g_{0}\|_{2,g}.\end{split}$
Then, from the definitions of scalar curvature rigid metrics of types I and II, if $g_{0}$ is a type I (resp. type II) scalar curvature rigid in the sense of Listing, then $F_{1,g_{0}}$ (resp. $F_{2,g_{0}}$) is an upper bound of $R_{min}$ as a functional on $\mathcal{M}(M)$. That is, it holds that $F_{1,g_{0}}(g)\geq R_{min}(g)$ (resp. $F_{2,g_{0}}(g)\geq R_{min}(g)$) for all $g\in\mathcal{M}(M)$.
Proof.
Suppose there is a metric $g\in\mathcal{M}(M)$ such that $R_{min}(g)=\min_{M}R_{g}>F_{1,g_{0}}(g)$. Then, from Definition 1.1, there is a positive constant $c>0$ such that $g=c\cdot g_{0}$. Hence, $R_{g}=c^{-1}R_{g_{0}}=c^{-1}F_{1,g_{0}}(g_{0})=F_{1,g_{0}}(g)$ on $M$. This contradicts our supposition $R_{min}(g)=\min_{M}R_{g}>F_{1,g_{0}}(g)$. The proof for the corresponding statement to $F_{2,g_{0}}$ is similar. ∎
[…]
As pointed out in [20, Section 3], when the Yamabe invariant, alias the sigma constant, $Y(M^{n})$ is non-positive, then
$\mathcal{M}\ni g\mapsto\mu(g):=Y(M^{n})\cdot\mathrm{Vol}(M,g)^{-2/n}\in\mathbb{R}$
is an upper bound of $R_{min}$ as a functional on $\mathcal{M}(M)$. Indeed, for a conformal class $C$ on $M$, if its Yamabe constant $Y(M,C)$ is non-positive, then
$Y(M,C)=\sup_{g\in C}R_{min}(g)\cdot\mathrm{Vol}(M,g)^{2/n}$
(see [13, Corollary 5.16]). Hence,
$Y(M)=\sup_{g\in\mathcal{M}(M)}R_{min}(g)\cdot\mathrm{Vol}(M,g)^{2/n}$
if $Y(M)\leq 0$. Moreover, if a smooth metric $g$ attains the equality, then it is a Yamabe metric (i.e., $Y(M,[g])=Y(M)$) and an Einstein metric (see [27]). Here, the Yamabe invariant $Y(M^{n})$ is defined as follows.
[…]
where the supremum is taken over all conformal classes on $M$.
3 Preliminaries
The following first variation formula of scalar curvature functional is well-known (see [3]).
Lemma 3.1.
$DR|_{\bar{g}}(h)=-\Delta_{\bar{g}}(tr_{\bar{g}}h)+\mathrm{div}_{\bar{g}}(\mathrm{div}_{\bar{g}}h)-\langle\mathrm{Ric}_{\bar{g}},h\rangle_{\bar{g}}.$ Here, $\Delta_{\bar{g}}f=\mathrm{tr}_{\bar{g}}\,\nabla_{\bar{g}}df$ is the non-positive Laplacian acting on the space of functions on $M$.
A more detailed calculation shows that if $g=\bar{g}+h$ for a metric $\bar{g}$ and a symmetric $(0,2)$-tensor $h$ with $\|h\|_{\bar{g}}<<1$, then
$R_{g}=\bar{R}+DR|_{\bar{g}}(h)+(g+h)^{-1}hg^{-1}hg^{-1}*\mathrm{Ric}_{\bar{g}}+g^{-1}*g^{-1}*g^{-1}*\bar{\nabla}h*\bar{\nabla}h,$ (15)
where the term $g^{-1}*g^{-1}*g^{-1}*\bar{\nabla}h*\bar{\nabla}h$ is a contraction of three copies of $g^{-1}$ (i.e., $g$ with raised indices) and two copies of ¯$\bar{\nabla}h=\bar{\nabla}g$. And, the term $(\bar{g}+h)^{-1}h\bar{g}^{-1}h\bar{g}^{-1}*\mathrm{Ric}_{\bar{g}}$ is the trace of $\mathrm{Ric}_{\bar{g}}$ with respect to $\left((\bar{g}+h)^{-1}h\bar{g}^{-1}h\bar{g}^{-1}\right)^{-1}$. Note that $\bar{g}+h$ is positive definite if $\|h\|_{\bar{g}}$ is small enough. All the proofs of these formulas are given in Section 7 below.
Take $h=u\cdot g$ for some smooth function $u\in C^{\infty}(M)$ on $M$. Then for any $s\in\mathbb{R}$,
$\begin{split}R_{g+th}&(g+sh)\\ &=R_{g}g+(DR|_{g}(h)g+R_{g}h)s\\ &~{}~{}~{}+\left(\frac{s^{2}u^{2}}{1+su}R_{g}+\frac{s^{2}}{4}(1-su^{-1})^{3}(2n-2)|\nabla u|^{2}\right)(g+sh)\\ &=R_{g}g-s(n-1)(\Delta_{g}u)g\\ &~{}~{}~{}+(1+su)\left(\frac{s^{2}u^{2}}{1+su}R_{g}+\frac{s^{2}}{4}(1-su^{-1})^{3}(2n-2)|\nabla u|^{2}\right)g.\end{split}$ (16)
Hence, if $M$ is a compact manifold with non-empty boundary $\partial M$, then we can take $u$ as the first eigenfunction of $\Delta_{g}$ and obtain the following.
Proposition 3.1.
Let $(M^{n},g)~{}(n\geq 2)$ be a compact Riemannian $n$-manifold with non-empty boundary $\partial M$. Let $u\in C^{\infty}(M)$ be the first eigenfunction of $\Delta_{g}$ with Dirichlet boundary condition. Then there is a small $s>0$ (resp. $s<0$) such that the metric $g_{s,u}:=(1+su)g$ satisfies that
$R_{g_{s,u}}\cdot g_{s,u}>R_{g}\cdot g~{}~{}(\mathrm{resp.}~{}R_{g_{s,u}}\cdot g_{s,u}<R_{g}\cdot g)$ (17)
at each point in the interior of $M$ and $g_{s,u}=g$ on $\partial M$.
If, moreover, $\frac{\partial u}{\partial\nu_{g}}$ is positive (resp. negative) everywhere on the boundary $\partial M$, then
$H_{g_{s,u}}>H_{g}~{}(\mathrm{resp.}~{}H_{g_{s,u}}<H_{g})~{}~{}\mathrm{on}~{}\partial M$ (18)
for sufficiently small $s>0$. Here, $\nu_{g}$ is the unit normal vector field on $\partial M$ of $g$.
Let $x=(x^{1},\cdots,x^{n},x^{n+1})$ be the Cartesian coordinate of $\mathbb{R}^{n+1}$ and $\mathbb{S}^{n}_{+}:=\{x\in\mathbb{R}^{n+1}\mid x^{n+1}\geq 0\}\subset\mathbb{R}^{n+1}$ the upper hemisphere. Then since the coordinate function $x^{n+1}$ is homogeneous function in $\mathbb{R}^{n+1}$ of degree one, its restriction to $\mathbb{S}^{n}_{+}$ is an eigenfunction of $\Delta_{\delta_{\mathbb{S}^{n}_{+}}}$, whose eigenvalue is $n$. Here, $\delta_{\mathbb{S}^{n}_{+}}$ is the restriction of the Euclidean metric $\delta$. Therefore $(\mathbb{S}^{n}_{+},g_{s,u}:=(1+sx^{n+1}|_{\mathbb{S}^{n}_{+}})\delta|_{\mathbb{S}^{n}_{+}})$ satisfies (17) and (18) for sufficiently small $s>0$. On the other hand, in order to construct a similar example of metric $h$ on $\mathbb{S}^{n}_{+}$ satisfying (18) and
$R_{h}>n(n-1)=R_{\delta|_{\mathbb{S}^{n}_{+}}}~{}\mathrm{on}~{}\mathbb{S}^{n}_{+}$
instead of (17), a more subtle deformation is needed (see [4, Theorem 4]).
Proposition 3.2.
Let $(M^{n},g)$ be a complete non-compact smooth Riemannian manifold. Assume that $\mathrm{Ric}_{g}\geq-K(n-1)$ for some $K\geq 0$ and $R_{g}>0$ on $M$. Moreover, assume that there is a smooth positive function $u\in C^{\infty}(M)$ satisfying
$-\Delta_{g}u=\lambda u$
for some positive constant $\lambda>0$. Then, for any $r>0$, there is a small $s>0$ (resp. $s<0$) depending only on $n,g,\lambda,K$ and $r$ such that $g_{s,u}:=(1+su)g$ is a smooth metirc in $B_{r}(p)$ and that
$R_{g_{s,u}}\cdot g_{s,u}>R_{g}\cdot g~{}~{}(\mathrm{resp.}~{}R_{g_{s,u}}\cdot g_{s,u}<R_{g}\cdot g)$
at each point of $B_{r}(p)$.
Proof.
First, we prove the case of $s>0.$ From the formula (16) and Li-Wang’s gradient estimate [18, Theorem 6.1], there is a constant $C>0$ depending on $n$ such that
$\begin{split}R_{g_{s,u}}\cdot g_{s,u}&\geq R_{g}\cdot g+s(n-1)(\lambda u)g+s^{2}(1+su)u^{-1}\left(\frac{R_{g}u^{3}}{1+su}-s^{3}C(r^{-2}+\lambda+K)\right)g\\ &\geq R_{g}\cdot g+s(n-1)(\lambda u)g\\ &~{}~{}~{}~{}+s^{2}(1+su)u^{-1}\min_{B_{r}(p)}R_{g}\left(\frac{u^{3}}{1+su}-\left(\min_{B_{r}(p)}R_{g}\right)^{-1}s^{3}C(r^{-2}+\lambda+K)\right)g\end{split}$
If $0\leq s\leq(\max_{B_{r}(p)}u)^{-1}$, then $(1+su)^{-1}\geq 1/2$. So, the desired assertion holds if
$0<s<\min\left\{\left(\frac{C^{-1}}{2}(r^{-2}+\lambda+K)^{-1}\left(\min_{B_{r}(p)}R_{g}\right)\min_{B_{r}(p)}u^{3}\right)^{1/3},\frac{1}{\max_{B_{r}(p)}u}\right\}.$
Next, we prove the case of $s<0$. If $0\geq s\geq-(\max_{B_{r}(p)}u)^{-1}$, then $1+su\geq 0$ and
$\begin{split}R_{g_{s,u}}\cdot g_{s,u}&\leq R_{g}\cdot g+s(n-1)(\lambda u)g\\ &~{}~{}~{}~{}+s^{2}u^{2}\left(R_{g}+(1+su)(1-su^{-1})^{3}C(r^{-2}+\lambda+K)\right)g.\end{split}$
If $0\geq s\geq\max\left\{-\frac{1}{\min_{B_{r}(p)}u},-\min_{B_{r}(p)}u\right\}=:s_{r,u}$,
$R_{g}+(1+su)(1-su^{-1})^{3}C(r^{-2}+\lambda+K)\leq\max_{B_{r}(p)}R_{g}+16C(r^{-2}+\lambda+K).$
Hence, the desired assertion holds if
$0>s>\max\left\{s_{r,u},-\frac{(n-1)\lambda}{\max_{B_{r}(p)}u\,(\max_{B_{r}(p)}R_{g}+16C(r^{-2}+\lambda+K))}\right\}$
4 Proofs of Main Theorems
From the observation in Section 3, on every closed $n$-manifold $M$, we cannot deform every metric on $M$ in the conformal direction so that the quantity $R_{g}\cdot g$ increases at each point of $M$. Indeed, the first order term (in terms of the parameter $s$) of the perturbed quantity $R_{(1+su)g}(1+su)g$ for a smooth function $u\in C^{\infty}(M)$ is
$-(n-1)(\Delta_{g}u)g.$
Thus, by the maximum principle, $\Delta_{g}u$ is sign-changing otherwise $u$ is constant on each connected component of $M$. Therefore, in order to increase the quantity $R_{g}\cdot g$ at each point on the closed manifold, we need to deform the given metric in a direction transverse to the conformal one. Let $M$ be a closed manifold and $\mathcal{M}$ the space of all (smooth) Riemannian metrics on $M$. From [9], for any metric $g\in\mathcal{M}$, the tangent space $T_{g}\mathcal{M}$ at $g$ is orthogonally decomposed as
$(C^{\infty}(M)\cdot g+\{\mathcal{L}_{X}g~{}|~{}X\in\Gamma(TM)\})\oplus TT,$
where the subspace $TT$ consists of tt-tensors which are trace-free and divergence-free (with respect to $g$) symmetric $(0,2)$-tensors. The traceless Ricci tensor
$\overset{\circ}{\mathrm{Ric}_{g}}:=\mathrm{Ric}_{g}-\frac{R_{g}}{n}g$
is orthogonal to the subspace $C^{\infty}(M)\cdot g$ (with respect to $L^{2}(M,g)$ inner product). Moreover, if $R_{g}$ is constant, then it is also a tt-tensor. Indeed, from the contracted second Bianchi identity:
$\delta\mathrm{Ric}_{g}=-\frac{1}{2}\nabla R_{g},$
and $R_{g}\equiv\mathrm{const}$, $\overset{\circ}{\mathrm{Ric}_{g}}$ is divergence-free. Since $\overset{\circ}{\mathrm{Ric}_{g}}$ is also trace-free, hence it is a tt-tensor. We are going to take this tt-tensor $\overset{\circ}{\mathrm{Ric}_{g}}$ as the variation $h$ in the following proof of our main theorems.
Proof of Theorem 1.3.
From the formula (15), we have for $10<s<<$ 1,
$\begin{split}&R_{g-s(\mathrm{Ric}_{g}-\frac{R_{g}}{n}g)}-R_{g}\cdot\|g\|^{2}_{1,\,g-s(\mathrm{Ric}_{g}-\frac{R_{g}}{n}g)}\\ &=R_{g}+\frac{n-1}{n}s\Delta_{g}R_{g}-\frac{R_{g}^{2}}{n}s\\ &~{}~{}~{}~{}+\frac{s}{2}\Delta_{g}R_{g}+s\|\mathrm{Ric}_{g}\|^{2}_{g}-\frac{R_{g}}{1-s\overline{\lambda}_{Ric}+\frac{R_{g}}{n}s}\\ &+g_{s}^{-1}\left(-s\left(\mathrm{Ric}_{g}-\frac{R_{g}}{n}g\right)\right)g^{-1}\left(-s\left(\mathrm{Ric}_{g}-\frac{R_{g}}{n}g\right)\right)g^{-1}*\mathrm{Ric}_{g}\\ &+Q_{s}(\nabla^{g}\mathrm{Ric}_{g}),\end{split}$
where $g_{s}:=g-s\left(\mathrm{Ric}_{g}-\frac{R_{g}}{n}g\right)$. We have used here that
  • $DR|_{g}(-\mathrm{Ric}_{g})=\frac{1}{2}\Delta_{g}R_{g}+\|\mathrm{Ric}_{g}\|_{g}^{2}$, and
  • $DR|_{g}\left(R_{g}\right)=-(n-1)(\Delta_{g}R_{g})g$.
The first one is known as the first variation of the scalar curvature functional along the Ricci flow (see [1, Corollary 4.20] for example). Here, $\overline{\lambda}_{Ric}(x)~{}(x\in M)$ is the largest eigenvalue of $\mathrm{Ric}_{g}(x)$ on $T_{x}M$, and $Q_{s}(\nabla^{g}\mathrm{Ric}_{g})=O(s^{2})\cdot Q(\nabla^{g}\mathrm{Ric}_{g})$, where $Q(\nabla^{g}\mathrm{Ric}_{g})$ is a contraction of three copies of $g_{s}^{-1}$ and two copies of $\nabla^{g}\mathrm{Ric}_{g}$.
$\begin{split}&\left(g-s\left(\mathrm{Ric}_{g}-\frac{R_{g}}{n}g\right)\right)^{-1}\\ &=\left(1+s\frac{R_{g}}{n}\right)^{-1}g^{-1}+s\left(1+s\frac{R_{g}}{n}\right)^{-2}g^{-1}\mathrm{Ric}_{g}g^{-1}+O(s^{2}).\end{split}$
Therefore,
$\begin{split}\left(g-s\left(\mathrm{Ric}_{g}-\frac{R_{g}}{n}g\right)\right)^{-1}&\left(-s\left(\mathrm{Ric}_{g}-\frac{R_{g}}{n}g\right)\right)g^{-1}\left(-s\left(\mathrm{Ric}_{g}-\frac{R_{g}}{n}g\right)\right)g^{-1}*\mathrm{Ric}_{g}\\ &=\frac{s^{2}\frac{R_{g}^{3}}{n^{2}}}{1+s\frac{R_{g}}{n}}+O(s^{2})\\ &=-R_{g}\left(1-s\frac{R_{g}}{n}\right)+\frac{R_{g}}{1+s\frac{R_{g}}{n}}+O(s^{2}).\end{split}$
Substituting this into the formula above, we have
$\begin{split}R_{g_{s}}&-R_{g}\cdot\|g\|^{2}_{1,g_{s}}\\ &=s\|\overset{\circ}{\mathrm{Ric}}_{g}\|^{2}_{g}+\frac{3n-2}{2n}\Delta_{g}R_{g}+s\frac{R^{2}_{g}}{n}+\frac{R_{g}}{1+s\frac{R_{g}}{n}}-\frac{R_{g}}{1-s\overline{\lambda}_{Ric}+\frac{R_{g}}{n}s}+O(s^{2})\\ &=s\|\overset{\circ}{\mathrm{Ric}}_{g}\|^{2}_{g}+\frac{3n-2}{2n}\Delta_{g}R_{g}+s\frac{R^{2}_{g}}{n}\\ &~{}~{}~{}~{}~{}-R_{g}\left(\frac{s\overline{\lambda}_{Ric}}{\left(1-s\overline{\lambda}_{Ric}+s\frac{R_{g}}{n}\right)\left(1+s\frac{R_{g}}{n}\right)}\right)+O(s^{2})\\ &=s\|\overset{\circ}{\mathrm{Ric}}_{g}\|^{2}_{g}+\frac{3n-2}{2n}\Delta_{g}R_{g}\\ &~{}~{}~{}~{}~{}-\frac{sR_{g}}{\left(1-s\overline{\lambda}_{Ric}+s\frac{R_{g}}{n}\right)\left(1+s\frac{R_{g}}{n}\right)}\left(\overline{\lambda}_{Ric}-\frac{R_{g}}{n}+O(s)\right)+O(s^{2})\end{split}$ (19)
Then the desired result for the case of $s>0$ follows directly from this formula. The claim for the case of $s<0$ can be proved similarly by replacing $\overline{\lambda}_{Ric}$ with $\underline{\lambda}_{Ric}$ in the above argument. Here $\underline{\lambda}_{Ric}(x)~{}(x\in M)$ is the smallest eigenvalue of $\mathrm{Ric}_{g}(x)$ on $T_{x}M$
Proof of Theorem 1.4.
Let $\lambda_{Ric,1}(x)\leq\cdots\leq\lambda_{Ric,n-1}(x)\leq\lambda_{Ric,n}(x)$ be the eigenvalues of $\mathrm{Ric}_{g}(x)$ on $T_{x}M~{}(x\in M)$. As in the proof of Theorem 1.3, we have that for all sufficiently small $10<|s|<<$ 1 (so that $(1+(R_{g}/n)s)g-s\mathrm{Ric}_{g}$ is positive definite on $M$),
if $R_{g}\leq 0$ and $s>0$,
$\begin{split}R_{g_{s}}&-R_{g}\cdot\|g\|_{2,g_{s}}\\ &=s\|\overset{\circ}{\mathrm{Ric}}_{g}\|^{2}_{g}+s\frac{R^{2}_{g}}{n}+\frac{R_{g}}{1+s\frac{R_{g}}{n}}+\frac{3n-2}{2n}\Delta_{g}R_{g}\\ &~{}~{}~{}-\frac{R_{g}}{\sqrt{\left(1+\frac{R_{g}}{n}s-s\lambda_{Ric,n}\right)\left(1+\frac{R_{g}}{n}s-s\lambda_{Ric,n-1}\right)}}+O(s^{2})\\ &\geq s\|\overset{\circ}{\mathrm{Ric}}_{g}\|^{2}_{g}+\frac{3n-2}{2n}\Delta_{g}R_{g}+s\frac{R^{2}_{g}}{n}+\frac{R_{g}}{1+s\frac{R_{g}}{n}}-\frac{R_{g}}{1-s\lambda_{Ric,n-1}+\frac{R_{g}}{n}s}+O(s^{2})\\ &=s\|\overset{\circ}{\mathrm{Ric}}_{g}\|^{2}_{g}+\frac{3n-2}{2n}\Delta_{g}R_{g}+s\frac{R^{2}_{g}}{n}\\ &~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}-R_{g}\left(\frac{s\lambda_{Ric,n-1}}{\left(1-s\lambda_{Ric,n-1}+s\frac{R_{g}}{n}\right)\left(1+s\frac{R_{g}}{n}\right)}\right)+O(s^{2})\\ &=s\|\overset{\circ}{\mathrm{Ric}}_{g}\|^{2}_{g}+\frac{3n-2}{2n}\Delta_{g}R_{g}\\ &~{}~{}~{}~{}-\frac{sR_{g}}{\left(1-s\lambda_{Ric,n-1}+s\frac{R_{g}}{n}\right)\left(1+s\frac{R_{g}}{n}\right)}\left(\lambda_{Ric,n-1}-\frac{R_{g}}{n}+O(s)\right)+O(s^{2}).\end{split}$
If $R_{g}\geq 0$ and $s>0$,
$\begin{split}R_{g_{s}}&-R_{g}\cdot\|g\|_{2,g_{s}}\\ &=s\|\overset{\circ}{\mathrm{Ric}}_{g}\|^{2}_{g}+s\frac{R^{2}_{g}}{n}+\frac{R_{g}}{1+s\frac{R_{g}}{n}}+\frac{3n-2}{2n}\Delta_{g}R_{g}\\ &~{}~{}~{}-\frac{R_{g}}{\sqrt{\left(1+\frac{R_{g}}{n}s-s\lambda_{Ric,n}\right)\left(1+\frac{R_{g}}{n}s-s\lambda_{Ric,n-1}\right)}}+O(s^{2})\\ &\geq s\|\overset{\circ}{\mathrm{Ric}}_{g}\|^{2}_{g}+\frac{3n-2}{2n}\Delta_{g}R_{g}+s\frac{R^{2}_{g}}{n}+\frac{R_{g}}{1+s\frac{R_{g}}{n}}-\frac{R_{g}}{1-s\lambda_{Ric,n}+\frac{R_{g}}{n}s}+O(s^{2})\\ &=s\|\overset{\circ}{\mathrm{Ric}}_{g}\|^{2}_{g}+\frac{3n-2}{2n}\Delta_{g}R_{g}+s\frac{R^{2}_{g}}{n}\\ &~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}-R_{g}\left(\frac{s\lambda_{Ric,n}}{\left(1-s\lambda_{Ric,n}+s\frac{R_{g}}{n}\right)\left(1+s\frac{R_{g}}{n}\right)}\right)+O(s^{2})\\ &=s\|\overset{\circ}{\mathrm{Ric}}_{g}\|^{2}_{g}+\frac{3n-2}{2n}\Delta_{g}R_{g}\\ &~{}~{}~{}~{}-\frac{sR_{g}}{\left(1-s\lambda_{Ric,n}+s\frac{R_{g}}{n}\right)\left(1+s\frac{R_{g}}{n}\right)}\left(\lambda_{Ric,n}-\frac{R_{g}}{n}+O(s)\right)+O(s^{2}).\end{split}$
Then, the assertion corresponding to $s>0$ follows directly from these estimates.
For the case of $s<0$, one can prove the desired assertion by replacing $\lambda_{Ric,n},\lambda_{Ric,n-1}$ with $\lambda_{Ric,2},\lambda_{Ric,1}$ respectively for each case: $R_{g}\leq$ and $\geq 0$ respectively. (Of course, the second inequality sign should be the opposite for each estimate in this case.) ∎
Proof of Theorem 1.5.
From the formula (15), we have
$\begin{split}&R_{g_{s}}\cdot g_{s}\\ &=\left(R_{g}+\frac{n-1}{n}s\Delta_{g}R_{g}-s\frac{R_{g}^{2}}{n}+\frac{s}{2}\Delta_{g}R_{g}+s\|\mathrm{Ric}_{g}\|^{2}_{g}+O(s^{2})\right)\left(g-s\left(\mathrm{Ric}_{g}-\frac{R_{g}}{n}g\right)\right)\\ &=R_{g}\cdot g+\left(\frac{n-1}{n}s\Delta_{g}R_{g}+\frac{s}{2}\Delta_{g}R_{g}+s\left(\|\overset{\circ}{\mathrm{Ric}_{g}}\|_{g}^{2}+\frac{R_{g}^{2}}{n}\right)\right)\cdot g-sR_{g}\cdot\mathrm{Ric}_{g}+O(s^{2})\end{split}$
on $TM$. Hence the desired assertion follows directly from this formula. ∎
Proof of Theorem 1.6.
As the proof of Theorem 1.5, we have
$\begin{split}&R^{2}_{g_{s}}\cdot g_{s}\\ &=\left(R_{g}+\frac{n-1}{n}s\Delta_{g}R_{g}-s\frac{R_{g}^{2}}{n}+\frac{s}{2}\Delta_{g}R_{g}+s\|\mathrm{Ric}_{g}\|^{2}_{g}+O(s^{2})\right)^{2}\left(g-s\left(\mathrm{Ric}_{g}-\frac{R_{g}}{n}g\right)\right)\\ &=R^{2}_{g}\cdot g+s\left(\frac{3n-2}{n}R_{g}\,\Delta_{g}R_{g}+2R_{g}\,\|\overset{\circ}{\mathrm{Ric}_{g}}\|_{g}^{2}+\frac{R_{g}^{3}}{n}\right)\cdot g-sR^{2}_{g}\cdot\mathrm{Ric}_{g}+O(s^{2})\end{split}$
on $\Lambda^{2}TM$. Hence the desired assertion follows directly from this formula. ∎
Proof of Theorem 1.7.
Let $h$ be a tt-tensor on $(M,g)$. Then, as discussed in [16, Remark 1.6], one can find a curve $g_{s}~{}(|s|<<1)$ lying in the space of all Riemannian metrics of constant scalar curvature with unit volume such that $\frac{d}{ds}g_{s}|_{s=0}=h$. By the Taylor expansion, we have that
$R_{g_{s}}=R_{g}+DR|_{g}(h)\cdot s+\frac{1}{2}\left(D^{2}R|_{g}(h,h)+DR|_{g}\left(h_{2}\right)\right)\cdot s^{2}+O(s^{3})$
on $M$. Here, $h_{2}:=\left.\frac{d^{2}}{d^{2}s}g_{s}\right|_{s=0}$. Since $g$ is an Einstein metric and $h$ is a tt-tensor, $DR|_{g}(h)=0$. Hence, for any $\varepsilon>0$ there is a small $s>0$ such that
$\begin{split}R_{g_{s}}-R_{g}\cdot&\|g\|^{2}_{1,g_{s}}\geq R_{g}+\frac{1}{2}\left(D^{2}R|_{g}(h,h)+DR|_{g}(h_{2})\right)\cdot s^{2}+O(s^{3})-\frac{R_{g}}{1+s(h_{min}\pm\varepsilon)}\\ &=R_{g}+\frac{1}{2}\left(D^{2}R|_{g}(h,h)+DR|_{g}(h_{2})\right)\cdot s^{2}\\ &~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}-R_{g}\left(1-s(h_{min}\pm\varepsilon)+s^{2}(h_{min}\pm\varepsilon)^{2}\right)+O(s^{3})\\ &=sR_{g}(h_{min}\pm\varepsilon)-\left(\frac{1}{4}\langle\Delta_{E}h,h\rangle_{L^{2}}-\frac{1}{2}DR|_{g}(h_{2})+R_{g}(h_{min}\pm\varepsilon)^{2}\right)s^{2}+O(s^{3}),\end{split}$
Here,
±$\pm\varepsilon:=\begin{cases}+\varepsilon&\mathrm{if}~{}R_{g}\leq 0~{}\mathrm{and}\\ -\varepsilon&\mathrm{if}~{}R_{g}\geq 0,\end{cases}$
$h_{min}(x):=\sup\{C\in\mathbb{R}\mid h_{x}(v,v)\geq C\cdot g_{x}(v,v)~{}\mathrm{for~{}all}~{}v\in T_{x}M\}\leq 0$ since $\mathrm{tr}_{g}h=0$, and we have used
$D^{2}R|_{g}(h,h)=-\frac{1}{2}\langle\Delta_{E}h,h\rangle_{L^{2}}$
since $g_{s}$ lies in the space of all Riemannian metrics of constant scalar curvature (see [16, Remark 1.6]). Therefore, if $R_{g}=0$ and $DR_{g}(h_{2})=0$, then $R_{g_{s}}>R_{g}\cdot\|g\|^{2}_{1,g_{s}}$ on $M$ for all sufficiently small $s>0$. If $R_{g}<0$ and $\max_{M}h_{min}<0$, then one can take $s>0$ small enough so that $h_{min}+\varepsilon<0$ on $M$. Therefore, $R_{g_{s}}>R_{g}\cdot\|g\|^{2}_{1,g_{s}}$ on $M$ for all sufficiently small $s>0$. Note that since $\mathrm{tr}_{g}h=0$, $\max_{M}h_{min}<0$ is equivalent to that $h$ is non-vanishing on $M$. ∎
Proof of Theorem 1.8.
As in the proof of Theorem 1.6, for any $\varepsilon>0$ there is a small $s>0$ such that
$\begin{split}&R_{g_{s}}-R_{g}\cdot\|g\|_{2,g_{s}}\geq R_{g}-\left(\frac{1}{4}\langle\Delta_{E}h,h\rangle_{L^{2}}-\frac{1}{2}DR|_{g}(h_{2})\right)\cdot s^{2}+O(s^{3})\\ &~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}-R_{g}\cdot\begin{cases}\frac{1}{\sqrt{(1+s(\lambda^{h}_{1}+\varepsilon))(1+s(\lambda^{h}_{2}+\varepsilon))}}&\mathrm{if}~{}R_{g}\leq 0\\ \frac{1}{\sqrt{(1+s(\lambda^{h}_{1}-\varepsilon))(1+s(\lambda^{h}_{2}-\varepsilon))}}&\mathrm{if}~{}R_{g}\geq 0\end{cases}\\ &\geq R_{g}-\left(\frac{1}{4}\langle\Delta_{E}h,h\rangle_{L^{2}}-\frac{1}{2}DR|_{g}(h_{2})\right)\cdot s^{2}+O(s^{3})+\begin{cases}\frac{-R_{g}}{1+s(\lambda^{h}_{2}+\varepsilon)}&\mathrm{if}~{}R_{g}\leq 0\\ \frac{-R_{g}}{1+s(\lambda^{h}_{1}-\varepsilon)}&\mathrm{if}~{}R_{g}\geq 0\end{cases}\\ &=R_{g}-\left(\frac{1}{4}\langle\Delta_{E}h,h\rangle_{L^{2}}-\frac{1}{2}DR|_{g}(h_{2})\right)\cdot s^{2}+O(s^{3})\\ &~{}~{}~{}~{}~{}~{}~{}~{}-R_{g}\cdot\begin{cases}(1-s(\lambda^{h}_{2}+\varepsilon)+s^{2}(\lambda^{h}_{2}+\varepsilon)^{2})&\mathrm{if}~{}R_{g}\leq 0\\ (1-s(\lambda^{h}_{1}-\varepsilon)+s^{2}(\lambda^{h}_{1}-\varepsilon)^{2})&\mathrm{if}~{}R_{g}\geq 0\end{cases}\\ &=\begin{cases}sR_{g}(\lambda^{h}_{2}+\varepsilon)-\left(\frac{1}{4}\langle\Delta_{E}h,h\rangle_{L^{2}}-\frac{1}{2}DR|_{g}(h_{2})+R_{g}(\lambda^{h}_{2}+\varepsilon)^{2}\right)s^{2}+O(s^{3})&\mathrm{if}~{}R_{g}\leq 0\\ sR_{g}(\lambda^{h}_{1}-\varepsilon)-\left(\frac{1}{4}\langle\Delta_{E}h,h\rangle_{L^{2}}-\frac{1}{2}DR|_{g}(h_{2})+R_{g}(\lambda^{h}_{1}-\varepsilon)^{2}\right)s^{2}+O(s^{3})&\mathrm{if}~{}R_{g}\geq 0.\end{cases}\end{split}$
Similarly, if $s<0$ with sufficiently small $|s|$, we have the same estimate by replacing $\lambda^{h}_{1},\varepsilon$ with $\lambda^{h}_{2},-\varepsilon$ respectively. Note that $\lambda^{h}_{1}\leq 0$ since $\mathrm{tr}_{g}h=0$, hence the desired assertions follow from these estimates. ∎
Remark 4.1.
From (19), there is a continuous function $F$ on $M$, which is determined by $M,n$ and $g$ such that one can show (2) of Theorem 1.3 on any compact subset $K\subset M$ provided that
$\inf\left\{s>0:~{}\|\overset{\circ}{\mathrm{Ric}}_{g}\|_{g}^{2}-R_{g}\left(\overline{\lambda}_{Ric}-\frac{R_{g}}{n}\right)>sF~{}\mathrm{on}~{}K\right\}>0.$ (20)
In particular, if the minimum of
$\|\overset{\circ}{\mathrm{Ric}}_{g}\|_{g}^{2}-R_{g}\left(\overline{\lambda}_{Ric}-\frac{R_{g}}{n}\right)$ (21)
on $K$ is positive, then the above condition (20) holds. For any compact set $K\subset\{x\in M\mid~{}\|\overset{\circ}{\mathrm{Ric}_{g}}\|_{g}(x)\neq 0\}$, the condition: $R_{g}\equiv\mathrm{const}\leq 0$ on $K$, especially implies (21). Similarly, there is a continuous symmetric $(0,2)$-tensor $h$ on $M$, which is determined by $M,n$ and $g$ such that one can show (7) of Theorem 1.5 on any compact subset $K\subset M$ provided that
$\inf\left\{s>0:~{}\left(\|\overset{\circ}{\mathrm{Ric}_{g}}\|_{g}^{2}+\frac{R^{2}_{g}}{n}\right)g-R_{g}\cdot\mathrm{Ric}_{g}>sh~{}\mathrm{on}~{}K\right\}>0.$ (22)
In particular, the condition: $R_{g}\equiv\mathrm{const}\leq 0$ on $K$ and $\min_{K}\underline{\lambda}_{Ric}\geq\frac{R_{g}}{n}$, especially implies (22).
5 Examples
5.1 Examples on Lie groups
We can construct examples of left invariant metrics on some three dimensional Lie groups that satisfy the assumptions of our main theorems in Section 1. Curvatures of left invariant metrics on Lie groups have been studied by Milnor in [26]. Especially, the case of three-dimensional unimodular is written in Section 4 of [26]. The following are examples of metrics on unimodular three-dimensional Lie groups, which satisfy the assumption of Theorem 1.3 or 1.5.
  • A left invariant metric on $SU(2)$ (which is homeomorphic to the unit 3-sphere $\mathbb{S}^{3}$), whose signature of Ricci quadratic form is (+,,)(+,-,-)( + , - , - ) (see also [12, Example 2]). More specifically, we consider the Berger sphere:
    $\begin{split}\mathbb{S}^{3}(1)\cong SU(2)&=\left\{A\in M(2,\mathbb{C})\mid~{}\det A=1,~{}A^{*}=-A\right\}\\ &=\left\{\begin{pmatrix}z&-\bar{w}\\ w&\bar{z}\end{pmatrix}\mid~{}(z,w)\in\mathbb{C}^{2},~{}|z|^{2}+|w|^{2}=1\right\},\end{split}$
    and set
    $X_{1}=\begin{pmatrix}\sqrt{-1}&0\\ 0&-\sqrt{-1}\end{pmatrix},~{}X_{2}=\begin{pmatrix}0&1\\ -1&0\end{pmatrix},~{}X_{3}=\begin{pmatrix}0&\sqrt{-1}\\ \sqrt{-1}&0\end{pmatrix}.$
    We define the left invariant metric $g_{s,t}~{}(1\leq s\leq t)$ on $SU(2)$ so that
    $g_{s,t}(X_{1},X_{1})=1,g_{s,t}(X_{2},X_{2})=s,g_{s,t}(X_{3},X_{3})=t,g_{s,t}(X_{i},X_{j})=0~{}(i\neq j).$
    Then, using the orthonormal frame
    $v_{1}:=X_{1},~{}v_{2}:=\frac{1}{\sqrt{s}}X_{2},~{}v_{3}:=\frac{1}{\sqrt{t}}X_{3},$
    one can compute the Ricci and scalar curvatures as follows.
    $\begin{split}\mathrm{Ric}_{g_{s,t}}(v_{1},v_{1})&=-\frac{1}{st}(-2+2t^{2}+2s^{2}-4st),\\ \mathrm{Ric}_{g_{s,t}}(v_{2},v_{2})&=-\frac{1}{st}(2+2t^{2}-2s^{2}-4t),\\ \mathrm{Ric}_{g_{s,t}}(v_{3},v_{3})&=-\frac{1}{st}(2-2t^{2}+2s^{2}-4s),\end{split}$
    and
    $R_{g_{s,t}}=\frac{2}{st}\{2(s+t+st)-(1+s^{2}+t^{2})\}.$
    For example, when $(s,t)=(1,4-\varepsilon/2)~{}(\varepsilon\leq 6)$,
    $\mathrm{Ric}_{g_{s,t}}(v_{i},v_{i})=-4+\varepsilon~{}(i=1,2),~{}\mathrm{Ric}_{g_{s,t}}(v_{3},v_{3})=8-\varepsilon,$
    and
    $\|\overset{\circ}{\mathrm{Ric}}_{g_{s,t}}\|_{g_{s,t}}^{2}=\frac{3}{2}\left(8-\frac{4}{3}\varepsilon\right)^{2},~{}R_{g_{s,t}}\left(\overline{\lambda}_{Ric}-\frac{R_{g_{s,t}}}{3}\right)=\varepsilon\left(8-\frac{4}{3}\varepsilon\right).$
    Hence, if $40<\varepsilon<$ 4, then $g_{s,t}$ has a positive constant scalar curvature and satisfies the assumptions of Theorems 1.3, 1.4, 1.5 and 1.6. Moreover, from [12, Example 2], if $\varepsilon\leq 2$ (resp. $\varepsilon<2$), then $g_{s,t}$ ia a (resp. unique) Yamabe metric with positive scalar curvature. On the other hand, if $\varepsilon\leq 0$, then $g_{s,t}$ has a non-positive constant scalar curvature and satisfies the assumptions of Theorems 1.3, 1.4 and 1.5 but does not one of Theorem 1.6.
  • Any left invariant metric on the Heisenberg group
    $\left\{\begin{bmatrix}1&*&*\\ 0&1&*\\ 0&0&1\end{bmatrix}\in\mathbb{M}(\mathbb{R},3)\right\},$
    whose signature of Ricci quadratic form is (+,,)(+,-,-)( + , - , - ).
  • A left invariant metric on $SL(2,\mathbb{R})$ or $E(1,1)$, whose signature of Ricci quadratic form can be either (+,,)(+,-,-)( + , - , - ) or $(0,0,-)$ depending on the choice of the left invariant metric.
For each metric with constant negative scalar curvature in the above example, from [14],
$\mathcal{C}_{1}:=\{g\in\mathcal{M}\mid~{}R_{g}=\mathrm{const},~{}\mathrm{Vol}(M,g)=1\}$
is a (infinite dimensional) manifold near such a metric (after normalizing it so that it has unit volume). Here, $\mathcal{M}$ is the space of all Riemannian metrics on each manifold $M$. Moreover, the condition that
$\|\overset{\circ}{\mathrm{Ric}_{g}}\|_{g}(x)\neq 0~{}\mathrm{for~{}all}~{}x\in M$
is an open condition with respect to the $C^{\infty}$-topology on $\mathcal{M}$. Therefore, all metrics in $\mathcal{C}_{1}$ sufficiently $C^{\infty}$-close to such a metric also satisfy the assumption of Theorem 1.3.
5.2 Examples on the total spaces of Riemannian submersions with totally geodesic fibres
  • ([25, Section 5]) Let $\pi:S^{4n+3}\rightarrow\mathbb{H}P^{n}$ be the Hopf fibration whose fibres are the standard unit $3$-sphere $\mathbb{S}^{3}$. We denote the scalar curvatures of the fibres, the base space and the total space as $R^{F},R^{B}$ and $R^{M}$, respectively. Then,
    $R^{F}=6,~{}R^{B}=16n(n+2),~{}R^{M}=(4n+3)(4n+2).$
    The Ricci curvature of the canonical variation $g_{t}~{}(t>0)$ was calculated in [25] as follows.
    $\begin{split}\mathrm{Ric}^{t}(U,V)&=\left(\frac{R^{F}}{t\,\mathrm{dim}F}+t\left(\frac{R^{M}}{\mathrm{dim}F+\mathrm{dim}B}-\frac{R^{F}}{\mathrm{dim}F}\right)\right)\,g_{t}(U,V),\\ \mathrm{Ric}^{t}(X,Y)&=\left(\frac{R^{B}}{t\,\mathrm{dim}B}+t\left(\frac{R^{M}}{\mathrm{dim}F+\mathrm{dim}B}-\frac{R^{B}}{\mathrm{dim}B}\right)\right)\,g_{t}(X,Y),\end{split}$
    where $U,V$ are vertical vectors and $X,Y$ are horizontal vectors. So, in this example,
    $\mathrm{Ric}^{t}(U,V)=\left(\frac{2}{t}+4nt\right)\,g_{t}(U,V),~{}\mathrm{Ric}^{t}(X,Y)=\left(4(n+2)-6t\right)\,g_{t}(X,Y)$
    and
    $R_{g_{t}}=\frac{6}{t}-12nt+16n(n+2).$
    Therefore,
    $\begin{split}\|\overset{\circ}{\mathrm{Ric}_{g_{t}}}\|^{2}_{g_{t}}&=3\left(\frac{2}{t}+4nt-\frac{1}{4n+3}\left(\frac{6}{t}-12nt+16n(n+2)\right)\right)^{2}\\ &+4n\left(4(n+2)-6t-\frac{1}{4n+3}\left(\frac{6}{t}-12nt+16n(n+2)\right)\right)^{2}\\ &=3\left(\frac{8n}{4n+3}\cdot\frac{1}{t}+\frac{4n(4n+6)}{4n+3}t-\frac{16n(n+2)}{4n+3}\right)^{2}\\ &+4n\left(\frac{6}{4n+3}\cdot\frac{1}{t}+\frac{12n+18}{4n+3}t-\frac{4(n+2)(4n-1)}{4n+3}\right)^{2}.\end{split}$
    Let $\lambda^{V}_{Ric_{t}}$ and $\lambda^{H}_{Ric_{t}}$ be the eigenvalues of $\mathrm{Ric}_{g_{t}}$ in the vertical and horizontal directions, respectively. One can observe that
    $\begin{cases}\lambda_{Ric_{t}}^{V}\geq\lambda_{Ric_{t}}^{V}&0<t\leq\frac{1}{2n+3}~{}\mathrm{or}~{}t\geq 1\\ \lambda_{Ric_{t}}^{H}\geq\lambda_{Ric_{t}}^{H}&\frac{1}{2n+3}\leq t\leq 1.\end{cases}$
    Matsuzawa [25] observed that the canonical variation $g_{t}$ is an Einstein metric on $\mathbb{S}^{4n+3}$ if and only if $t=1$ or $t=\frac{1}{2n+3}$. Then, we have
    $\begin{split}&R_{g_{t}}\left(\lambda_{Ric_{t}}^{V}-\frac{R_{g_{t}}}{\mathrm{dim}M}\right)\\ &=\left(\frac{6}{t}-12nt+16n(n+2)\right)\left(\frac{8n}{4n+3}\cdot\frac{1}{t}+\frac{4n(4n+6)}{4n+3}t-\frac{16n(n+2)}{4n+3}\right),\\ &R_{g_{t}}\left(\lambda_{Ric_{t}}^{H}-\frac{R_{g_{t}}}{\mathrm{dim}M}\right)\\ &=\left(\frac{6}{t}-12nt+16n(n+2)\right)\left(\frac{6}{4n+3}\cdot\frac{1}{t}+\frac{12n+18}{4n+3}t-\frac{4(n+2)(4n-1)}{4n+3}\right).\end{split}$
    Hence, if $10<t<<$ 1 or $t>>1$, then
    $\|\overset{\circ}{\mathrm{Ric}_{g_{t}}}\|_{g_{t}}^{2}\cdot g_{t}>R_{g_{t}}\left(\mathrm{Ric}_{g_{t}}-\frac{R_{g_{t}}}{\mathrm{dim}M}\cdot g_{t}\right).$
    Therefore, $(\mathbb{S}^{4n+3},g_{t})$ satisfies the assumption of Theorems 1.3, 1.4 and 1.5 for all $10<t<<$ 1 and $t>>1$. Moreover, since $R_{g_{t}}>0$ on $M$, $g_{t}$ also satisfies the assumption of Theorem 1.6 for all $10<t<<$ 1 and $t>>1$.
  • ([25, Section 5]) Consider the fibration $\mathbb{C}P^{2n+1}\rightarrow\mathbb{H}P^{n}$ whose fibres are $S^{2}(4)$ with constant sectional curvature $4$. Then,
    $R^{F}=8,~{}R^{B}=16n(n+2),~{}R^{M}=(4n+4)(4n+2).$
    Let $g_{t}$ be the canonical variation of this Riemannian submersion. Then, as in the previous example, one can calculate as
    $\mathrm{Ric}^{t}(U,V)=\left(\frac{4}{t}+4nt\right)\,g_{t}(U,V),~{}\mathrm{Ric}^{t}(X,Y)=\left(4(n+2)-4t\right)\,g_{t}(X,Y)$
    and
    $R_{g_{t}}=\frac{8}{t}-8nt+16n(n+2).$
    Therefore,
    $\begin{split}\|\overset{\circ}{\mathrm{Ric}_{g_{t}}}\|^{2}_{g_{t}}&=2\left(\frac{4}{t}+4nt-\frac{1}{4n+2}\left(\frac{8}{t}-8nt+16n(n+2)\right)\right)^{2}\\ &+4n\left(4(n+2)-4t-\frac{1}{4n+2}\left(\frac{8}{t}-8nt+16n(n+2)\right)\right)^{2}\\ &=2\left(\frac{8n}{2n+1}\cdot\frac{1}{t}+\frac{2n(4n+4)}{2n+1}t-\frac{8n(n+2)}{2n+1}\right)^{2}\\ &+4n\left(\frac{4}{2n+1}\cdot\frac{1}{t}+\frac{4(n+1)}{2n+1}t-\frac{4(n+2)}{2n+1}\right)^{2}.\end{split}$
    Let $\lambda^{V}_{Ric_{t}}$ and $\lambda^{H}_{Ric_{t}}$ be the eigenvalues of $\mathrm{Ric}_{g_{t}}$ in the vertical and horizontal directions, respectively. One can observe that
    $\begin{cases}\lambda_{Ric_{t}}^{V}\geq\lambda_{Ric_{t}}^{V}&0<t\leq\frac{1}{n+1}~{}\mathrm{or}~{}t\geq 1\\ \lambda_{Ric_{t}}^{H}\geq\lambda_{Ric_{t}}^{H}&\frac{1}{n+1}\leq t\leq 1.\end{cases}$
    Matsuzawa [25] observed that the canonical variation $g_{t}$ is an Einstein metric on $\mathbb{S}^{4n+3}$ if and only if $t=1$ or $t=\frac{1}{n+1}$. Then, we have
    $\begin{split}&R_{g_{t}}\left(\lambda_{Ric_{t}}^{V}-\frac{R_{g_{t}}}{\mathrm{dim}M}\right)\\ &=\left(\frac{8}{t}-8nt+16n(n+2)\right)\left(\frac{8n}{2n+1}\cdot\frac{1}{t}+\frac{2n(4n+4)}{2n+1}t-\frac{8n(n+2)}{2n+1}\right),\\ &R_{g_{t}}\left(\lambda_{Ric_{t}}^{H}-\frac{R_{g_{t}}}{\mathrm{dim}M}\right)\\ &=\left(\frac{8}{t}-8nt+16n(n+2)\right)\left(\frac{4}{2n+1}\cdot\frac{1}{t}+\frac{4(n+1)}{2n+1}t-\frac{4(n+2)}{2n+1}\right).\end{split}$
    Hence, if $10<t<<$ 1 or $t>>1$, then
    $\|\overset{\circ}{\mathrm{Ric}_{g_{t}}}\|_{g_{t}}^{2}\cdot g_{t}>R_{g_{t}}\left(\mathrm{Ric}_{g_{t}}-\frac{R_{g_{t}}}{\mathrm{dim}M}\cdot g_{t}\right).$
Therefore, $(\mathbb{C}P^{2n+1},g_{t})$ satisfies the assumption of Theorems 1.3, 1.4 and 1.5 for all $10<t<<$ 1 and $t>>1$. Moreover, since $R_{g_{t}}>0$ on $M$, $g_{t}$ also satisfies the assumption of Theorem 1.6 for all $10<t<<$ 1 and $t>>1$.
6 Conclusion
6.1 Other rigidity results
Bär [2] recently proved the following rigidity result.
Theorem 6.1 ([2, Main Theorem]).
Let $(M,g)$ be a connected closed Riemannian spin manifold of dimension $\geq 2$ and $D$ the Dirac operator acting on spinor fields of $M$. Then
$\lambda_{1}(D^{2})\leq\frac{n^{2}}{4\,\mathrm{Rad}_{\mathbb{S}^{n}}(M,d_{g})}.$ (23)
Equality holds in (23) if and only if $(M,g)$ is isometric to $(\mathbb{S}^{n}(R),g_{std})$ with $R=\mathrm{Rad}_{\mathbb{S}^{n}}(M,d_{g})$.
Here, $\mathrm{Rad}_{\mathbb{S}^{n}}(M,d_{g})$ is the hyperspherical radius of $M$, which is defined as the supremum of all numbers $R>0$ such that there exists a Lipschitz map $f:(M,f_{g})\rightarrow(\mathbb{S}^{n},d_{g_{std}})$ with Lipschitz constant $\mathrm{Lip}(f)\leq 1/R$ and $\mathrm{deg}(f)\neq 0$.
As mentioned in subsection 3.1 in [2], if $\min_{M}R_{g}>0$, then
$\mathrm{Rad}_{\mathbb{S}^{n}}(M,d_{g})^{2}\leq\frac{n(n-1)}{\min_{M}R_{g}},$
and equality holds if and only if $(M,g)$ is isometric to $(\mathbb{S}^{n},g_{std})$. This especially implies Llarull’s rigidity theorem ([21, 6, 17]) [^1] See the remarks immediately following the Theorem 1 in [2].. Note that for any Lipschitz map $f:M\rightarrow\mathbb{S}^{n}$ with $\mathrm{deg}(f)\neq 0$,
$\mathrm{Lip}(f)\geq\frac{1}{\mathrm{Rad}_{\mathbb{S}^{n}}(M,d_{g})}.$
Hence, for any Lipschitz map $f:M\rightarrow\mathbb{S}^{n}$ with $\mathrm{deg}(f)\neq 0$,
$\frac{R_{g}}{\mathrm{Lip}(f)^{2}}\geq\frac{R_{g_{std}}\circ f}{\mathrm{Lip}(\mathrm{id}_{\mathbb{S}^{n}})^{2}}=n(n-1)~{}~{}\mathrm{on}~{}M$ (24)
implies $f:(M,d_{g})\rightarrow(\mathbb{S}^{n},g_{std})$ is an isometry. Here, $\mathrm{id}_{\mathbb{S}^{n}}:\mathbb{S}^{n}\rightarrow\mathbb{S}^{n}$ denotes the identity map, hence $\mathrm{Lip}(\mathrm{id}_{\mathbb{S}^{n}})=1$.
Motivated by this, we call a Riemanniam manifold $(M_{0},g_{0})$ extremal in the sense of (24) if
$\frac{R_{g}}{\mathrm{Lip}(f)^{2}}\geq R_{g_{0}}\circ f~{}~{}\mathrm{on}~{}M$
for any Lipschitz map $f:M\rightarrow M_{0}$ with $\mathrm{deg}(f)\neq 0$ implies that $f:(M,d_{g})\rightarrow(M_{0},g_{0})$ is an isometry.
Question 6.1.
What kinds of properties does extremal metric in the sense of (24) have? Can we find sufficient conditions for a metric not to be extremal in this sense, as in our main theorems?
Similarly, from [2, Theorem 4], for any Lipschitz map $f:M\rightarrow\mathbb{S}^{n}$ with $\mathrm{deg}(f)\neq 0$,
$\frac{Y(M,[g])}{\mathrm{Lip}(f)^{2}\mathrm{Vol}(M,g)^{2/n}}\geq\frac{Y(\mathbb{S}^{n},[g_{std}])}{\mathrm{Lip}(\mathrm{id}_{\mathbb{S}^{n}})^{2}\mathrm{Vol}(\mathbb{S}^{n},g_{std})^{2/n}}$ (25)
implies that $(M,g)$ is isometric to $(\mathbb{S}^{n},g_{std})$. Here, $Y(M,[g])$ is the Yamabe constant of the conformal class $[g]$ on $M$ (see Remark 2.1 above).
Motivated by this, we call a Riemanniam manifold $(M_{0},g_{0})$ extremal in the sense of (25) if
$\frac{Y(M,[g])}{\mathrm{Lip}(f)^{2}\mathrm{Vol}(M,g)^{2/n}}\geq\frac{Y(M_{0},[g_{0}])}{\mathrm{Vol}(M_{0},g_{0})^{2/n}}$
for any Lipschitz map $f:M\rightarrow M_{0}$ with $\mathrm{deg}(f)\neq 0$ implies that $(M,g)$ is isometric to $(M_{0},g_{0})$.
Question 6.2.
What kinds of properties does extremal metric in the sense of (25) have? Can we find sufficient conditions for a metric not to be extremal in this sense, as in our main theorems?
6.2 For singular metrics
Question 6.3.
Let $M$ be a compact smooth manifold and $\Sigma\subset M$ is a closed subset. Can we find a sufficient condition for a metric not to be scalar curvature rigid with a given “boundary condition” associated with $\Sigma$. For example,
  • $\Sigma=\partial M$ and the “boundary condition” is something involving the mean curvature of $\partial M$, or
  • $\Sigma$ is an arbitrary closed subset and consider the set $\Sigma$ as the set of singular points of the metric in some sense. Here, the “boundary condition” is the decay of metric near the singular set $\Sigma$.
Next, we mention that the scalar minimum functional $R_{min}$ (Section 2.1) and a generalized definition of scalar curvature bounded below. Let $M^{n}$ be a smooth closed $n$-manifold and $\kappa\in\mathbb{R}$ a constant. Let $C^{0}_{met}(M,\kappa)$ denote the $C^{0}$-completion of $C^{2}$-metrics whose scalar curvature is bounded below by $\kappa$ in the conventional sense. That is, a $C^{0}$-metric $g$ is in $C^{0}_{met}(M,\kappa)$ if and only if there exists a sequences of $C^{2}$-metrics $g_{i}$ on $M$ such that $g_{i}$ converges uniformly to $g$ and satisfies $R(g_{i})\geq\kappa$. From the observation after Remark 1.10 of [5] and Theorem 1.7 of the same paper, one can observe that a $C^{0}$-metric $g$ is in $C^{0}_{met}(M,\kappa)$ if and only if
$\limsup_{C^{2}\ni h\rightarrow g,C^{0}}R_{min}(h)\geq\kappa.$
Here, “$\limsup_{C^{2}\ni h\rightarrow g,C^{0}}R_{min}(h)\geq\kappa$” means that
$\lim_{\delta\rightarrow 0}\left(\sup_{h\in\mathcal{M}^{2},~{}\|h-g\|_{C^{0}(M,g)}<\delta}R_{min}(h)\right)\geq\kappa,$
where $\mathcal{M}^{2}$ denotes the set of all $C^{2}$-metrics on $M$. Note that if
$\limsup_{C^{2}\ni h\rightarrow g,C^{0}}R_{min}(h):=\alpha<+\infty,$
then this is equivalent to that for any $\varepsilon>0$ there is $\delta>0$ such that
$R_{min}(h)<\alpha+\varepsilon$
for all $h\in\mathcal{M}^{2}$ with $\|h-g\|_{C^{0}(M,g)}<\delta$ and there exists a sequence of $C^{2}$-metrics $h_{i}$ on $M$ such that $R_{min}(h_{i})\rightarrow\alpha$ as $i\rightarrow\infty$.
Consider the following particular situation. Let $M$ be a closed manifold with non-positive Yamabe invariant $Y(M)\leq 0$. Then, as mentioned in Remark 2.1,
$Y(M)=\sup_{h\in\mathcal{M}}R_{min}(h)\cdot\mathrm{Vol}(M,g)^{2/n}.$
Assume that there is a sequence of $C^{2}$-metrics $g_{i}$ on $M$ such that $g_{i}\overset{C^{0}}{\longrightarrow}g$ and
$R_{min}(g_{i})\cdot\mathrm{Vol}(M,g_{i})^{2/n}\rightarrow Y(M)\,(\leq 0)~{}~{}\mathrm{as}~{}i\rightarrow\infty.$ (26)
Then, $R_{min}(g_{i})\rightarrow Y(M)\cdot\mathrm{Vol}(M,g)^{-2/n}$ as $i\rightarrow\infty$ and hence
$\limsup_{C^{2}\ni h\rightarrow g,C^{0}}R_{min}(h)\geq\lim_{i\rightarrow\infty}R_{min}(g_{i})=Y(M)\cdot\mathrm{Vol}(M,g)^{-2/n}.$
Therefore, $g\in C^{0}_{met}\left(M,Y(M)\cdot\mathrm{Vol}(M,g)^{-2/n}\right)$. Of course, a typical example of $(g_{i})$ that satisfies only (26) is a sequence of solutions to the Yamabe problem on each conformal class $[g_{i}]$.
Related to this observation, it is interesting to explore some relations between variational properties of $R_{min}$ in a topology weaker than $C^{2}$ and singular Yamabe metrics or other extremal metrics.
6.3 About our assumptions
Question 6.4.
In our theorems 1.3, 1.4, 1.5 and 1.6, can we weaken the assumption that “$\|\overset{\circ}{\mathrm{Ric}_{g}}\|_{g}(x)\neq 0$ for all $x\in M$” to that “the metric $g$ is not an Einstein metric”?
Question 6.5.
Does every closed manifold $M$ of dimension $n\geq 3$ admit a metric $g$ with non-positive constant scalar curvature and $\|\overset{\circ}{\mathrm{Ric}_{g}}\|_{g}(x)\neq 0$ for all $x\in M$?
Several examples of manifolds that admit no Einstein metric are known. On the other hand, every closed manifold of dimension $\geq 3$ admits a metric with constant negative scalar curvature. Therefore, such a manifold of dimension $\geq 3$ always admits a metric of non-Einstein negative constant scalar curvature. Moreover, Matsuo [24, Corollary 2] proved that there exists a non-Ricci-flat scalar-flat metric on every closed manifold of dimension $\geq 3$ with positive Yamabe invariant. However, one cannot distinguish whether or not the norm of the traceless Ricci tensor of each such metric has a positive lower bound on the whole manifold.
Question 6.6.
Are there any examples that satisfy the assumptions of Theorem 1.7 or 1.8 (especially the scalar non-flat case)?
7 Appendix
Let $g$ and $\bar{g}$ be two Riemannian metrics on a $n$-manifold. Set the difference between the Levi-Civita connections of $g$ and $\bar{g}$ as
$W:=\nabla-\bar{\nabla}.$
Then $W$ is a tensor (unlike $\Gamma$). With respect to a local frame $e_{1},\cdots,e_{n}$, we can write the components of $W$ via
$(\nabla_{i}-\bar{\nabla}_{i})(e_{j})=W^{k}_{ij}e_{k}.$
First, direct computations deduce the following two propositions.
Proposition 7.1.
In a local coordinates,
$W^{k}_{ij}=\frac{1}{2}g^{kl}(\bar{\nabla}_{i}g_{lj}+\bar{\nabla}_{j}g_{il}-\bar{\nabla}_{l}g_{ij}).$
Here, ¯$\bar{\nabla}_{i}g_{jk}$ denotes the expression of $\nabla^{\bar{g}}g$ in terms of the local coordinates.
Proof.
¯$\begin{split}\bar{\nabla}_{i}g_{lj}&=\partial_{i}g_{lj}-g_{pj}\bar{\Gamma}^{p}_{il}-g_{lp}\bar{\Gamma}^{p}_{ij}\\ \bar{\nabla}_{j}g_{il}&=\partial_{j}g_{il}-g_{pl}\bar{\Gamma}^{p}_{ji}-g_{ip}\bar{\Gamma}^{p}_{jl}\\ -\bar{\nabla}_{l}g_{ij}&=-\partial_{l}g_{ij}+g_{pj}\bar{\Gamma}^{p}_{il}+g_{ip}\bar{\Gamma}^{p}_{lj}.\end{split}$
Taking the sum of these both sides, we get
$g^{kl}(\bar{\nabla}_{i}g_{lj}+\bar{\nabla}_{j}g_{il}-\bar{\nabla}_{l}g_{ij})=2\Gamma^{k}_{ij}-\delta^{k}_{p}\bar{\Gamma}^{p}_{ij}-\delta^{k}_{p}\bar{\Gamma}^{p}_{ji}=2\Gamma^{k}_{ij}-2\bar{\Gamma}^{k}_{ij}=2W^{k}_{ij}.$
Proposition 7.2.
In a local coordinates,
$R_{ij}=\bar{R}_{ij}+\bar{\nabla}_{k}W^{k}_{ij}-\bar{\nabla}_{i}W^{k}_{kj}+W^{p}_{ij}W^{k}_{kp}-W^{p}_{kj}W^{k}_{ip}.$ (27)
Here, $R_{ij}$ and $\bar{R}_{ij}$ denote the expressions of the Ricci tensors of $g$ and $\bar{g}$ respectively, in terms of the local coordinates.
Proof.
$\begin{split}W^{p}_{ij}W^{k}_{kp}&=(\Gamma^{p}_{ij}-\bar{\Gamma}^{p}_{ij})(\Gamma^{k}_{kp}-\bar{\Gamma}^{k}_{kp})=\Gamma^{p}_{ij}\Gamma^{k}_{kp}-\underset{(1)}{\underline{\bar{\Gamma}^{p}_{ij}\Gamma^{k}_{kp}}}-\underset{(2)}{\underline{\Gamma^{p}_{ij}\bar{\Gamma}^{k}_{kp}}}+\underset{(3)}{\underline{\bar{\Gamma}^{p}_{ij}\bar{\Gamma}^{k}_{kp}}}\\ -W^{p}_{kj}W^{k}_{ip}&=-(\Gamma^{p}_{kj}-\bar{\Gamma}^{p}_{kj})(\Gamma^{k}_{ip}-\bar{\Gamma}^{k}_{ip})=-\Gamma^{p}_{kj}\Gamma^{k}_{ip}+\underset{(4)}{\underline{\bar{\Gamma}^{p}_{kj}\Gamma^{k}_{ip}}}+\underset{(6)}{\underline{\Gamma^{p}_{kj}\bar{\Gamma}^{k}_{ip}}}-\underset{(5)}{\underline{\bar{\Gamma}^{p}_{kj}\bar{\Gamma}^{k}_{ip}}}\\ \bar{\nabla}_{k}W^{k}_{ij}&=\bar{\nabla}_{k}(\Gamma^{k}_{ij}-\bar{\Gamma}^{k}_{ij})\\ &=\partial_{k}\Gamma^{k}_{ij}-\partial_{k}\bar{\Gamma}^{k}_{ij}+\underset{(2)}{\underline{\Gamma^{p}_{ij}\bar{\Gamma}^{k}_{kp}}}-\underset{(7)}{\underline{\Gamma^{k}_{pj}\bar{\Gamma}^{p}_{ik}}}-\underset{(4)}{\underline{\Gamma^{k}_{ip}\bar{\Gamma}^{p}_{kj}}}-\bar{\Gamma}^{p}_{ij}\bar{\Gamma}^{k}_{kp}+\underset{(8)}{\underline{\bar{\Gamma}^{k}_{pj}\bar{\Gamma}^{p}_{ik}}}+\underset{(5)}{\underline{\bar{\Gamma}^{k}_{ip}\bar{\Gamma}^{p}_{kj}}}\\ -\bar{\nabla}_{i}W^{k}_{kj}&=-\bar{\nabla}_{i}(\Gamma^{k}_{kj}-\bar{\Gamma}^{k}_{kj})\\ &=-\partial_{i}\Gamma^{k}_{kj}+\partial_{i}\bar{\Gamma}^{k}_{kj}-\underset{(6)}{\underline{\Gamma^{p}_{kj}\bar{\Gamma}^{k}_{ip}}}+\underset{(7)}{\underline{\Gamma^{k}_{pj}\bar{\Gamma}^{p}_{ki}}}+\underset{(1)}{\underline{\Gamma^{k}_{kp}\bar{\Gamma}^{p}_{ij}}}+\bar{\Gamma}^{p}_{kj}\bar{\Gamma}^{k}_{ip}-\underset{(8)}{\underline{\bar{\Gamma}^{k}_{pj}\bar{\Gamma}^{p}_{ki}}}-\underset{(3)}{\underline{\bar{\Gamma}^{k}_{kp}\bar{\Gamma}^{p}_{ij}}}.\end{split}$
Therefore, we have (where terms with the same number cancel each other)
¯$\begin{split}\bar{\nabla}_{k}W^{k}_{ij}&-\bar{\nabla}_{i}W^{k}_{kj}+W^{p}_{ij}W^{k}_{kp}-W^{p}_{kj}W^{k}_{ip}\\ &=\partial_{k}\Gamma^{k}_{ij}-\partial_{i}\Gamma^{k}_{kj}+\Gamma^{p}_{ij}\Gamma^{k}_{kp}-\Gamma^{p}_{kj}\Gamma^{k}_{ip}\\ &~{}~{}~{}~{}~{}-(\partial_{k}\bar{\Gamma}^{k}_{ij}-\partial_{i}\bar{\Gamma}^{k}_{kj}+\bar{\Gamma}^{p}_{ij}\bar{\Gamma}^{k}_{kp}-\bar{\Gamma}^{p}_{kj}\bar{\Gamma}^{k}_{ip})\\ &=R_{ij}-\bar{R}_{ij}.\end{split}$
Proposition 7.3.
$DR|_{\bar{g}}(h)=-\Delta_{\bar{g}}(tr_{\bar{g}}h)+\mathrm{div}_{\bar{g}}(\mathrm{div}_{\bar{g}}h)-\langle\mathrm{Ric}_{\bar{g}},h\rangle_{\bar{g}}.$
Proof.
From Proposition 7.2 ($g=g_{t}:=\bar{g}+th~{}(|t|<<1)$), we have
$R_{g_{t}}=(\bar{g}^{ij}-th^{ij})\bar{R}_{ij}+(\bar{g}^{ij}-th^{ij})(\bar{\nabla}_{k}W^{k}_{ij}-\bar{\nabla}_{j}W^{k}_{ki})+\mathrm{other~{}terms}.$
Now, since $W|_{t=0}=0$, the “other terms” is vanishing when $t=0$. From Proposition 7.1, we have
$W^{k}_{ij}=\frac{1}{2}(\bar{g}^{kl}-t\bar{h}^{kl})\left(\bar{\nabla}_{i}(\bar{g}_{lj}+th_{lj})+\bar{\nabla}_{j}(\bar{g}_{il}+th_{il})-\bar{\nabla}_{l}(\bar{g}_{ij}+th_{ij})\right).$
Since ¯$\bar{\nabla}\bar{g}=0$, we have
$\frac{d}{dt}W^{k}_{ij}|_{t=0}=\frac{1}{2}\bar{g}^{kl}(\bar{\nabla}_{i}h_{lj}+\bar{\nabla}_{j}h_{il}-\bar{\nabla}_{l}h_{ij}).$
Summing these up, we have
$\begin{split}\frac{d}{dt}R_{g_{t}}|_{t=0}&=-\langle\mathrm{Ric}_{\bar{g}},h\rangle+\bar{g}^{ij}\left(\bar{\nabla}_{k}\frac{d}{dt}W^{k}_{ij}|_{t=0}-\frac{d}{dt}\bar{\nabla}_{j}W^{k}_{ki}|_{t=0}\right)\\ &=-\langle\mathrm{Ric}_{\bar{g}},h\rangle+\frac{1}{2}\bar{g}^{ij}\left(\bar{\nabla}_{k}\bar{g}^{kl}(\bar{\nabla}_{i}h_{lj}+\bar{\nabla}_{j}h_{il}-\bar{\nabla}_{l}h_{ij})\right)\\ &-\frac{1}{2}\bar{g}^{ij}\left(\bar{\nabla}_{j}\bar{g}^{kl}(\bar{\nabla}_{k}h_{li}+\bar{\nabla}_{i}h_{kl}-\bar{\nabla}_{l}h_{ki})\right)\\ &=-\langle\mathrm{Ric}_{\bar{g}},h\rangle+\frac{1}{2}\bar{g}^{ij}\bar{g}^{kl}\bar{\nabla}_{k}(\underset{(1)}{\underline{\bar{\nabla}_{i}h_{lj}}}+\bar{\nabla}_{j}h_{il}-\underset{(2)}{\underline{\bar{\nabla}_{l}h_{ij}}})\\ &-\frac{1}{2}\bar{g}^{ij}\bar{g}^{kl}\bar{\nabla}_{j}(\bar{\nabla}_{k}h_{li}+\underset{(2)}{\underline{\bar{\nabla}_{i}h_{kl}}}-\underset{(1)}{\underline{\bar{\nabla}_{l}h_{ki}}})\\ &=-\langle\mathrm{Ric}_{\bar{g}},h\rangle-\underset{(2)}{\underline{\bar{g}^{kl}\bar{\nabla}_{k}\bar{\nabla}_{l}(\bar{g}^{ij}h_{ij})}}+\underset{(1)}{\underline{\bar{g}^{ij}\bar{g}^{kl}\bar{\nabla}_{k}\bar{\nabla}_{i}h_{lj}}}\\ &=-\langle\mathrm{Ric}_{\bar{g}},h\rangle-\bar{\Delta}(\mathrm{tr}_{\bar{g}}h)+\mathrm{div}_{\bar{g}}(\mathrm{div}_{\bar{g}}h).\end{split}$
(The terms with the same number have canceled each other.)
Here we have used
  • $W|_{t=0}=0$ in the 1st equality,
  • ¯$\bar{\nabla}\bar{g}=0$ in the 3rd and 4th equality,
  • In the 4th equality, the other term is vanishing.
Moreover, a more detailed calculation shows that if $g=\bar{g}+h~{}(\|h\|_{\bar{g}}<<1)$,
$R_{g}=\bar{R}+DR|_{\bar{g}}(h)+(\bar{g}+h)^{-1}h\bar{g}^{-1}h\bar{g}^{-1}*\mathrm{Ric}_{\bar{g}}+g^{-1}*g^{-1}*g^{-1}*\bar{\nabla}h*\bar{\nabla}h,$
where the term $g^{-1}*g^{-1}*g^{-1}*\bar{\nabla}h*\bar{\nabla}h$ is a contraction of three copies of $g^{-1}$ (i.e., $g$ with raised indices) and two copies of ¯$\bar{\nabla}h=\bar{\nabla}g$. And, the term $(\bar{g}+h)^{-1}h\bar{g}^{-1}h\bar{g}^{-1}*\mathrm{Ric}_{\bar{g}}$ is the trace of $\mathrm{Ric}_{\bar{g}}$ with respect to $\left((\bar{g}+h)^{-1}h\bar{g}^{-1}h\bar{g}^{-1}\right)^{-1}$. Note that $\bar{g}+h$ is positive definite if $\|h\|_{\bar{g}}$ is small enough. Indeed, this formula follows by taking the both sides of (27) in Proposition 7.2 with respect to $g=\bar{g}+h$ and using
$(\bar{g}+h)^{-1}=\bar{g}^{-1}-\bar{g}^{-1}h\bar{g}^{-1}+(\bar{g}+h)^{-1}h\bar{g}^{-1}h\bar{g}^{-1}.$
The term $g^{-1}*g^{-1}*g^{-1}*\bar{\nabla}h*\bar{\nabla}h$ comes from the “other terms” in the proof of Proposition 7.3.
Acknowledgements
The author was supported by JSPS KAKENHI Grant Number 24KJ0153.